OpenStudy (anonymous):

solve the equation for 3sin2θ - 2cos2θ = 2 for 0≤θ≤2π using the substitution t= tanθ

OpenStudy (anonymous):

\[3*\frac{ 2 \tan \theta }{ 1+\tan ^2\theta }-2*\frac{ 1-\tan ^2\theta }{ 1+\tan ^2 \theta }=2\] \[6 \tan \theta-2+2\tan ^2\theta =2+2\tan ^2\theta \] \[6 \tan \theta=2+2=4\] \[\tan \theta=\frac{ 3 }{ 2 }\] \[\theta=\tan^{-1} \frac{ 3 }{ 2 },\pi+\tan^{-1} \frac{ 3 }{ 2 }\]